OpenStudy (anonymous):

Solve: x(3x+1)(2x-5)=0

OpenStudy (tfht):

(3x^2+x)(2x-5)= (6x^3)+(2x^2)+(15x^2)+(-5x) (6x^3)+(17x^2)+(-5x)=0 x((6x^2)+(17x)-5 so then set everything to 0

OpenStudy (tfht):

wait nevermind... im lost

OpenStudy (tfht):

o set everything equal to 5 so x=5, x=the square root of 5/6 and x=5/17. I think